# Integral Springer Theorem for Quadratic Lattices under Base Change of Odd Degree

### v1

@@ 5.2 Proof of the main theorem
Suppose that $[E:F]$ is odd. Then, for each $\mathfrak{p}\in\mathrm{Spm}(R)$, there exists a $\mathfrak{q}\in\mathrm{Spm}(S)$ lying over $\mathfrak{p}$ such that $[E_{\mathfrak{q}}:F_{\mathfrak{p}}]$ is odd.
@@
When $E/F$ is separable, this follows from the fundamental equality $[E:F]=\sum_{\mathfrak{q}|\mathfrak{p}}[E_{\mathfrak{q}}:F_{\mathfrak{p}}].$ The case with $E/F$ purely inseparable of prime degree has been discussed in Remark 5.6. The general case follows easily from a combination of the above two special cases. ∎
@@
If $\widetilde{N}\hookrightarrow\mathrm{spn}^{+}(\widetilde{M})$ over $S$, then $N\hookrightarrow\mathrm{spn}^{+}(M)$ over $R$.
Similarly, if $\widetilde{N}\hookrightarrow\mathrm{spn}(\widetilde{M})$ over $S$, then $N\hookrightarrow\mathrm{spn}(M)$ over $R$.
@@
First assume that $\widetilde{N}\hookrightarrow\mathrm{spn}^{+}(\widetilde{M})$. Then by definition, there is an $S$-lattice $L\in\mathrm{spn}^{+}(\widetilde{M})\subseteq\mathrm{gen}(\widetilde{M})$ such that $\widetilde{N}$ is contained in $L$. Thus, for every $\mathfrak{q}\in\mathrm{Spm}(S)$, $S_{\mathfrak{q}}\otimes_{S}\widetilde{N}=S_{\mathfrak{q}}\otimes_{R}N$ is contained in $S_{\mathfrak{q}}\otimes_{S}L\cong S_{\mathfrak{q}}\otimes_{S}\widetilde{M}=S_{\mathfrak{q}}\otimes_{R}M$. For every $\mathfrak{p}\in\mathrm{Spm}(R)$, we can deduce from Lemma 5.7 and Theorem 2.2 that $N_{\mathfrak{p}}=R_{\mathfrak{p}}\otimes_{R}N$ can be embedded in $M_{\mathfrak{p}}=R_{\mathfrak{p}}\otimes_{R}M$ over $R_{\mathfrak{p}}$. So, by [28, 102:5], $N$ is embedded in some lattice $P\in\mathrm{gen}(M)$. Without loss of generality, we may assume $N\subseteq P$.
@@
Let $[\mathrm{spn}^{+}(M)]$ be the canonical image of $\mathrm{spn}^{+}(M)$ in the quotient
@@
We need to show that $[\mathrm{spn}^{+}(M)]=0$ in the above quotient group.
@@
where the horizontal maps are induced by the adelic spinor norm maps and the vertical maps are induced by the natural inclusions $\mathrm{O}^{+}(V)\hookrightarrow\mathrm{O}^{+}(V_{E})$ and $\mathbf{I}_{F}\hookrightarrow\mathbf{I}_{E}$. It is clear from the definition that the map $\theta_{F}$ in the diagram is injective. The assumption $\widetilde{N}\hookrightarrow\mathrm{spn}^{+}(\widetilde{M})$ means that $[\mathrm{spn}^{+}(M)]$ lies in the kernel of $\phi$. So it remains to prove the injectivity of the map $\psi$ in the above diagram.
@@
To prove the second assertion, assume that $\widetilde{N}\hookrightarrow\mathrm{spn}(\widetilde{M})$. In other words, $\widetilde{N}\subseteq\sigma\underline{\rho}(\widetilde{M})$ for some $\sigma\in\mathrm{O}(V_{E})$ and $\underline{\rho}\in\mathrm{O}_{\mathbf{A}}^{\prime}(V_{E})$. We may assume $\det(\sigma)=-1$. In this case choose any $\gamma\in\mathrm{O}(V)$ with $\det(\gamma)=-1$, then
$\widetilde{\gamma N}=\gamma\widetilde{N}\subseteq\gamma\sigma\underline{\rho}(\widetilde{M})\quad\text{with}\quad\gamma\sigma\underline{\rho}(\widetilde{M})\in\mathrm{spn}^{+}(\widetilde{M}).$
@@
If $\widetilde{N}$ is contained in some lattice in $\mathrm{cls}(\widetilde{M})$ (resp. in $\mathrm{cls}^{+}(\widetilde{M})$) over $S$, then $N$ is contained in some lattice in $\mathrm{cls}(M)$ (resp. in $\mathrm{cls}^{+}(M)$) over $R$.
@@
At this point the theorem follows from a combination of Corollary 5.9 with Remark 4.9, keeping in mind that for lattices $N$ and $M$ in a common quadratic space, $N$ can be embedded into $M$ if and only if $N$ is contained in some lattice in $\mathrm{cls}(M)$. ∎
@@ 5.3 Examples
The result of Theorem 1.2 can fail when $V=F\otimes_{R}M$ is 2-dimensional. To see this, consider the quadratic extension of $\mathbb{F}_{3}(t)$ given by $F=\mathbb{F}_{3}(t)(\sqrt{t^{3}-t^{2}+1})$ and let $R$ be the integral closure of $\mathbb{F}_{3}[t]$ inside $F$. Since the class number of $R$ is 5 by the table given in [1, p.232], there is a non-principal ideal $\mathfrak{a}$ of $R$ such that $\mathfrak{a}^{5}=rR$ for some $r\in R$ with $5]{r}\not\in F$. The quadratic lattice $M$ over $R$ defined by
@@
Indeed, the completion of $F$ at the infinite place (which corresponds to the $s$-adic valuation) is $F_{\infty}=\mathbb{F}_{q}(\!(s)\!)$. By Hensel’s lemma, the polynomial $f$ admits a factorization $f=gh$, where $g,\,h$ are monic polynomials with coefficients in $R^{\prime}=\mathbb{F}_{q}[\![s]\!]$ such that

### v2

@@ 5.2 Proof of the main theorem
Suppose that $[E:F]$ is odd. Then, for each $\mathfrak{p}\in\mathrm{Spm}(R)$, there exists a $\mathfrak{q}\in\mathrm{Spm}(S)$ lying over $\mathfrak{p}$ such that $E_{\mathfrak{q}}:F_{\mathfrak{p}}]$ is odd.
@@
When $E/F$ is separable, this follows from the fundamental equality $E:F]=\sum_{\mathfrak{q}|\mathfrak{p}}[E_{\mathfrak{q}}:F_{\mathfrak{p}}].$ The case with $E/F$ purely inseparable of prime degree has been discussed in Remark 5.6. The general case follows easily from a combination of the above two special cases. ∎
@@
[…]
@@
[…]
@@
Let $\mathrm{spn}^{+}(M)]$ be the canonical image of $\mathrm{spn}^{+}(M)$ in the quotient
@@
We need to show that $\mathrm{spn}^{+}(M)]=0$ in the above quotient group.
@@
[…]
@@
[…]
@@
[…]
@@
At this point the theorem follows from a combination of Corollary 5.9 with Remark 4.9, keeping in mind that for lattices $N$ and $M$ in a common quadratic space, NN can be embedded into $M$ if and only if $N$ is contained in some lattice in $\mathrm{cls}(M)$. ∎
@@ 5.3 Examples
The result of Theorem 1.2 can fail when $V=F\otimes_{R}M$ is 2-dimensional. To see this, consider the quadratic extension of $\mathbb{F}_{3}(t)$ given by $F=\mathbb{F}_{3}(t)(\sqrt{t^{3}-t^{2}+1})$ and let $R$ be the integral closure of $\mathbb{F}_{3}[t]$ inside $F$. Since the class number of $R$ is 5 by the table given in [1, p.232], there is a non-principal ideal $\mathfrak{a}$ of $R$ such that $\mathfrak{a}^{5}=rR$ for some $r\in R$ with $\sqrt[5]{r}\not\in F$. The quadratic lattice $M$ over $R$ defined by
@@
Indeed, the completion of $F$ at the infinite place (which corresponds to the $s$-adic valuation) is $F_{\infty}=\mathbb{F}_{q}(\!(s)\!)$. By Hensel’s lemma, the polynomial $f$ admits a factorization $f=gh$, where g,hg,\,h are monic polynomials with coefficients in $R^{\prime}=\mathbb{F}_{q}[\![s]\!]$ such that
@@
Acknowledgements
The first author was supported by the National Natural Science Foundation of China (grant no. 12171223). The second author was partially supported by the short-term visiting program for doctoral students at Southern University of Science and Technology. The third author was supported by the National Key R&D Program of China No. 2023YFA1009702 and the National Natural Science Foundation of China (grant no. 11631009).